The Hillcrest Garden Club wants to double the area of its rectangular rose bed. Strips of the same width will be added to one end and one side to form a rectangle. If the bed is now 17.5m by 12.2m, what are the dimensions of the new bed?
Answer by Theo(13342)

the area of the original rectangle is 12.2 * 17.5 = 213.5 square meters.

the new rectangle will have a length of 12.2 + x and a width of 17.5 + x.

the new area will be 2 *( 231.5 = 427 square meters.

you get (12.2 + x) * (17.5 + x) = 427.

solve this quadratic equation to get x = 5.9832066662876.

your new width will be 18.18320667.

your new length will be 23.48320667.

your new area will be 427 square meters.

a picture of what your new rectangle will look like is shown below:

the original rectangle is ABCD.

the new rectangle is EFGD.

$$$